OpenStudy (anonymous):

"solving exponential equations and inequalities" -solve each equation- ((1/2))^(x-3)=16^(3x+1) help PLEASE

OpenStudy (mertsj):

Rewrite the two sides with the same base.

OpenStudy (mertsj):

\[\frac{1}{2}^{x-3}=(2^{-1})^{x-3}=2^{-x+3}\]

OpenStudy (mertsj):

\[16^{3x+1}=(2^4)^{3x+1}=2^{12x+4}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

wait then what will x equalt to?

OpenStudy (mertsj):

If the bases are the same and the expressions are equal then the exponents must be equal.

OpenStudy (mertsj):

Set them equal and solve the resulting equation.